# HARMONIC: Cognitive and Control Collaboration in Human-Robotic Teams

### v2

@@ Abstract
This paper introduces HARMONIC, a cognitive-robotic architecture that integrates the OntoAgent cognitive framework with general-purpose robot control systems applied to human-robot teaming (HRT). We also present a cognitive strategy for robots that incorporates metacognition, natural language communication, and explainability capabilities required for collaborative partnerships in HRT. Through simulation experiments involving a joint search task performed by a heterogeneous team of a UGV, a drone, and a human operator, we demonstrate the system’s ability to coordinate actions between robots with heterogeneous capabilities, adapt to complex scenarios, and facilitate natural human-robot communication. Evaluation results show that robots using the OntoAgent architecture within the HARMONIC framework can reason about plans, goals, and team member attitudes while providing clear explanations for their decisions, which are essential prerequisites for realistic human-robot teaming.
@@ I Introduction
The proliferation of robots across various domains has given rise to a growing number of human-robot teams, where robots with various capabilities are employed by human operators to achieve complex mission goals [1]. Current state-of-the-art robotic systems demonstrate remarkable tactical proficiency, from the sub-millimeter precision of surgical robots performing intricate surgical procedures [2] to the dynamic multi-terrain locomotion of quadrupedal [3] and humanoid platforms exhibiting sophisticated whole-body coordination [4].
@@
These systems also excel at integrating multi-modal sensory information through sensor fusion, simultaneously localizing themselves within complex environments, recognizing objects, tracking moving targets with increasing reliability, in adaptive policy learning and multi-robot planning and collaboration [5, 6].
@@
Despite these significant advancements in tactical execution, most state-of-the-art robots are limited to short-horizon planning and execution and lack the strategic capabilities necessary for long-horizon planning, reasoning, and true collaborative partnership with humans in complex decision-making scenarios [7]. Importantly, effective human-robot teaming requires more than just tactical proficiency; to function as trusted partners within human-robot teams, they must exhibit a range of cognitive and metacognitive capabilities, including:
  • developing and enhancing an understanding of team tasks and team organization;
@@
  • developing and enhancing specific capabilities, involving the responsibilities, preferences, and actions of self and other team members [8];
@@
  • maintaining and using an episodic memory of past activities, as when leveraging past experiences to make informed predictions about future needs and actions;
@@
  • interpreting language, visual perception, and haptic perception in terms of an underlying formal model of the world (ontology); [9]
@@
  • communicating with teammates using interpreted, meaningful natural language.
@@
Figure 1: An overview of the framework, showing the Strategic and Tactical components representing the high-level planning (System 2) and low-level execution (System 1), respectively.
@@
Long-horizon functioning is within the scope of the field of cognitive architectures where much progress has been made in supporting complex agent team interactions, managing attention, setting goals, and conducting sophisticated planning, often incorporating diverse multi-agent planning strategies. Still, while cognitive architectures address key limitations in current robotic systems, they do not traditionally concentrate on the dynamic world of embodied agents. This has resulted in a notable disconnect between cognitive architectures and practical robotic control systems. Bridging this divide leverages the strengths of both these areas of research and development.
@@
The HARMONIC framework presented in this paper directly addresses this challenge through the use of a dual-layer architecture that combines strategic cognitive capabilities with tactical robotic control. This integration not only enables long-horizon planning but also enhances the scope of the robot’s reasoning capabilities, adds metacognitive awareness and social intelligence, which in turn supports human-level explanation that serves as a foundation for robots to function as trusted teammates [10], rather than mere tools, within human-robot teams.
The cognitive architecture that we incorporate into HARMONIC as its strategic layer is OntoAgent [11, 12, 9]. It is responsible for the aforementioned cognitive architecture capabilities, thus facilitating complex team interactions, attention management, goal-setting, sophisticated planning (including the incorporation of diverse multi-agent planning strategies), and addressing unexpected challenges in ways that humans can readily interpret.
Additionally, OntoAgent integrates both verbal and visual communication, significantly enhancing skill transfer and knowledge sharing within human-robot teams, grounding, and supporting meaning-oriented natural language communication, enabling clear causal explanations for the robots’ actions and decisions. These capabilities greatly enhance transparency and explainability [10], establishing crucial foundations for trust in human-robot collaborations [9].
@@
We present our approach to human-robot teaming using the example of a team comprised of two robots and a human executing a joint task. We demonstrate the robots’ knowledge of team organization as well as, their ability to (a) dynamically negotiate tasks and allocate responsibilities, (b) select and modify plans, and (c) adapt language communication to both the situation and the needs of their interlocutors. We implement these robotic capabilities using the novel HARMONIC architecture that operates in parallel at two levels – the strategic cognitive level, which supports operations requiring reasoning, and the tactical control level, which supports skill-based, ”automatic,” reflexive operation of robots.
Specifically, the key contributions of this paper are:
@@
  • We introduce HARMONIC, a cognitive-robotic architecture, which enables a flexible integration of OntoAgent [11], a cognitive architecture with low-level robot planning and control using Behavior Trees [13].
@@
  • Results of simulation experiments on a search task carried out by a robotic team, which showcases the human-robot team’s interactions and ability to handle complex, real-world scenarios.
Section II briefly describes the state of the art and background. Section III introduces the HARMONIC framework and the implementation of its strategic and tactical levels. Section IV-B describes the task, the team configuration, and the development of distributed plans, and Section IV presents the evaluation.
@@ II-A Cognitive-Robotic Architectures
There are relatively few cognitive architectures explicitly designed or effectively adapted for robotic applications due to the significant challenges involved in combining cognitive-level decision-making with motor control. Among these architectures, Soar [14], ACT-R [15], and DIARC [16] are prominent, each contributing distinct methods and philosophies toward cognitive autonomy in robotics.
@@
Finally, DIARC is explicitly designed for cognitive robotics with distributed modules [16, 17], affective processes, and advanced natural language capabilities suitable for dynamic human-robot interactions. However, it struggles with concurrent process management, lacks clear architectural layering, unified ontology, explainability, long-term episodic memory, explicit metacognition, and formalized safety mechanisms, thus limiting its robustness and adaptability.
In summary, most architectures require extensive customization from both cognitive and control perspectives, leading to limited real-world applications and practical testing.
@@ II-B Multi-Agent Planning (MAP) and Communication
Multi-Agent Planning (MAP) in Heterogeneous Multi-Robot (HMR) teams includes a variety of centralized and decentralized strategies, each presenting distinct trade-offs. Centralized planning achieves global optimality but encounters scalability challenges with larger teams, while decentralized methods improve scalability and adaptability but may result in suboptimal solutions [18], as highlighted by the exponential complexity identified by Brafman et al. [19]. Consequently, hybrid MAP approaches that blend centralized hierarchical planning with decentralized execution [20], dynamically switch between plans [5], or integrate centralized conflict resolution with decentralized planning [21] have gained traction. Advanced implementations of these hybrid strategies often incorporate auction-based task allocation systems and distributed planners [22, 23].
Regardless of the planning approach, effective intra-team communication remains integral for seamless coordination and execution within HMR teams. Such communication involves distinct protocols for human-robot and robot-robot interactions [24]. A unified, human-centric communication approach within a common cognitive architecture is advocated as essential by Lemon et al. [25] and Natarajan et al. [7]. Such architectures emulate human-like communication, reasoning, and team dynamics, building upon the notion of common ground introduced by Clark et al. [26] and further developed by Klein et al. [27]. Subsequent advances in dialogue processing systems [28] and grounding techniques [29, 30], together with insights from prior work on team organization [31] and natural language communication [9], guide current implementations of cognitive processes of OntoAgent in HARMONIC.
@@ II-C The OntoAgent Cognitive Architecture
OntoAgent is a content-centric cognitive architecture, designed to support the development of social intelligent agents through computational cognitive modeling [12, 9, 11]. This approach emphasizes the need to acquire, maintain, and dynamically expand large-scale knowledge bases, which are essential for an agent’s perception interpretation, reasoning, and action. The architecture’s memory structure is divided into three main components: a Situation Model (SM) that contains currently active concept instances; a Long-Term Semantic Memory (LTS) that stores knowledge about instances of events and objects, and an Episodic Memory (LTE) that stores knowledge about instances of events and objects. OntoAgent supports goal-directed behavior through a goal agenda and a prioritizer that selects which goals to pursue. It typically uses stored plans associated with goals but can also engage in reasoning from first principles when necessary.
The architecture is built on a service-based infrastructure, with key services including perception interpretation, attention management, reasoning, and action rendering. OntoAgent can integrate both native services and imported external capabilities, such as robotic vision or test-to-speech modules. This flexibility allows it to support multiple perception channels, including language understanding, visual perception, and simulated interoception (Chapter 8 of [12]).
@@
II-D Behavior Trees
Behavior Trees (BTs), originally developed for controlling non-player characters in video games, have found significant applications in robotics and AI [13]. BTs are directed trees that provide a flexible and intuitive way to design control actions and define task-planning hierarchies for agents. The structure of a BT includes control nodes (selectors, sequencers, and parallel nodes) that manage execution flow among child and execution nodes (action and condition nodes) that form the leaf nodes performing specific tasks. Control nodes process “ticks” at a set frequency, and conventionally execute from left to right, depending on the control node involved. This modular design enables easy modification and scaling of behaviors, making BTs particularly suitable for complex hierarchical task planning. In the current HARMONIC system, robot-specific BTs are integrated at the tactical layer to coordinate robotic control.
@@ III The HARMONIC Framework
HARMONIC (Human-AI Robotic Team Member Operating with Natural Intelligence and Communication), shown in Fig. 1, is a dual-control cognitive-robotic architecture that integrates strategic, cognitive-level decision-making with tactical, skill-level robot control. It builds upon and extends the concept of hybrid control systems and architectures, as discussed in a comprehensive review by Dennis et al. [32]. It also represents an advancement over the type 2 integration approach employed in DIARC [16, 17]. Unlike DIARC, where the strategic layer is embedded as a subsystem within the tactical layer to enable concurrent and dynamic operations, HARMONIC introduces a more sophisticated integration of these components.
@@
The tactical layer is responsible for attention management, lower-level robot planning and control, sensor input processing, and executing motor actions based on high-level commands received from the strategic layer. The attention manager processes perception data, translating and prioritizing sensory information for use by the strategic component. Additionally, the tactical layer employs specialized controllers, algorithms, and models to convert abstract commands from OntoAgent in the strategic layer into precise, actionable robot operations.
For example, a command such as PICK(key, at-position1), from the OntoAgent to ”pick up a key on the ground,” triggers a well-defined sequence of tasks and actions including object identification, positional assessment, trajectory computation, and actuation, in the tactical layer. Furthermore, the tactical layer manages reactive behaviors, including collision avoidance, through dedicated controllers.
The tactical layer of HARMONIC is built around the Behavior Trees (BTs) [13] which enable effective reactive control and support the system’s safety and operational requirements in dynamic environments. The tactical layer may also include advanced control policies for whole-body compliant control [33, 13] with motion planning, path planning [34] etc. In addition to these, BTs can also be used for representing skills and low-level plans that can also be shared between the robots [35, 36].
In the context of HARMONIC, state manager (a blackboard) keeps track of condition and state variables on the tactical layer [37]. As shown in Fig. 2, this allows for efficient querying and updating of the system’s state during operation based on the sensory inputs and action commands received from the strategic layer.
Sensory inputs are continuously written into the State Manager (SM), where the attention module filters and packages this perception data for the OntoAgent in the strategic layer. The high-level action commands sent by the OntoAgent through Interface APIs are unpacked for updating the state variables in the SM. While sensor data and physical state tracking operate in real time, the strategic layer runs asynchronously. Real-time control is achieved through the data available in the SM that is accessed by the LiveBT in parallel, which chooses policies and actions stored in the Action Manager. This architecture enables HARMONIC to maintain real-time behavior while supporting higher-level planning capabilities.
@@
Action commands from the strategic layer are translated into corresponding state and control variables via action data communication APIs. These APIs update the placeholder variables within the State Manager, ensuring accurate and timely control of the tactical layer operations.
The architecture’s design aligns with Kahneman’s dual-system approach [38], with the OntoAgent in the strategic layer implementing System 2 (slow, deliberative reasoning) and the BTs in the tactical layer implementing System 1 (fast, intuitive responses). This structure allows for dynamic scheduling and adaptation, enabling real-time adjustments to priorities and actions. It effectively handles computational delays, contingencies, safety concerns, and resource optimization through a combination of low-level planning and reactive algorithms.
@@ IV Task Design
We evaluated the HARMONIC framework’s teaming strategy using a simulated search task set in an apartment environment, as illustrated in Fig. 4.6. The scenario involves an HMR team consisting of an unmanned ground vehicle (UGV), a drone, and a human team member. The objective is to find a set of lost keys. The human initiates the task by remotely communicating with the robotic team members.
The simulation environment, developed using the Unity game engine, incorporates abstracted versions of robot behaviors. This approach allowed us to focus on key aspects of the HARMONIC framework’s performance while simplifying the implementation of complex robot actions that are not central to our current work.
@@ IV-A Heterogenous Robot Team Configuration
The simulation evaluation involved two robots: an Unmanned Ground Vehicle (UGV) and a drone, both functionally and structurally heterogeneous, offering complementary capabilities and perspectives [39]. Each robot is equipped with sensors for localization, collision detection, and object recognition, facilitating coordinated and efficient operations.
The UGV’s design (see Fig. 3.a) enables it to access ground-level areas and maneuver under obstacles and surfaces. The drone, on the other hand, effectively complements the UGV by performing aerial search operations. The drone’s (see Fig. 3.b) vertical mobility and sensing capabilities allow it to scan environments extensively from above, navigating over obstacles and accessing areas difficult for the UGV. Since the robots communicate in natural language, their dialogue and actions are comprehensible to humans, and humans can intervene when necessary.
Both robots operate decentralized instances of tactical and strategic modules, ensuring efficient and autonomous team planning and control. Currently, the UGV acts as the team leader at the task level, responsible for managing interactions with the human operator and coordinating with the drone. The leader assignment is presently randomized; however, future work aims to develop and evaluate heuristic methods for dynamic team hierarchies. This advancement will optimize leader selection based on contextual factors and the individual capabilities of each agent relevant to specific tasks.

### v3

@@ Abstract
This paper describes HARMONIC, a cognitive-robotic architecture that integrates the OntoAgent cognitive framework with general-purpose robot control systems applied to human-robot teaming (HRT). HARMONIC incorporates metacognition, meaningful natural language communication, and explainability capabilities required for developing mutual trust in HRT. Through simulation experiments involving a joint search task performed by a heterogeneous team of two HARMONIC-based robots and a human operator, we demonstrate heterogeneous robots that coordinate their actions, adapt to complex scenarios, and engage in natural human-robot communication. Evaluation results show that HARMONIC-based robots can reason about plans, goals, and team member attitudes while providing clear explanations for their decisions, which are essential requirements for realistic human-robot teaming.
@@ I Introduction
Current state-of-the-art robotic systems demonstrate remarkable tactical proficiency, from the sub-millimeter precision of surgical robots performing surgical procedures [2] to the dynamic multi-terrain locomotion of quadrupedal [3] and humanoid platforms exhibiting sophisticated whole-body coordination [4]. These systems also excel at integrating multimodal sensory information through sensor fusion, simultaneously localizing themselves within complex environments, recognizing objects, tracking moving targets with increasing reliability, in adaptive policy learning and multi-robot planning and collaboration [5, 6].
@@
Despite these significant advancements in tactical execution, most state-of-the-art robots are limited to short-horizon planning and execution and lack the strategic capabilities necessary for long-horizon planning, reasoning, and true collaborative partnership with humans in complex decision-making scenarios [7]. Importantly, effective human-robot teaming [1] requires more than just tactical proficiency; to function as trusted partners within human-robot teams, they must exhibit a range of cognitive and metacognitive capabilities, including:
  • developing and refining an understanding of team tasks and organization;
@@
  • building specific capabilities related to the responsibilities, preferences, and actions of self and other team members [8];
@@
  • maintaining and using episodic memory to leverage past experiences for predicting future needs and actions;
@@
  • interpreting language, visual, and haptic perception through an underlying formal model of the world (ontology) [9];
@@
  • communicating with teammates using meaningful, interpreted natural language.
@@
Figure 1: An overview of the HARMONIC framework, showing the Strategic and Tactical components representing the high-level planning (System 2) and low-level execution (System 1), respectively.
@@
Long-horizon functioning is within the scope of the field of cognitive systems, where much progress has been made on the tasks of complex agent team interactions, managing attention, goal selection, and plan execution. Still, while cognitive systems address key limitations in current robotic systems, they do not traditionally concentrate on the dynamic world of embodied agents. This has resulted in a disconnect between cognitive systems and practical robotic control systems. Our work integrates and leverages the strengths of both these areas of research and development.
@@
The HARMONIC framework uses a dual-layer architecture that integrates strategic cognitive capabilities with tactical robotic control. This integration not only enables long-horizon planning but also enhances the scope of the robot’s reasoning capabilities, adds metacognitive awareness and social intelligence, which in turn supports human-level explanation that serves as a foundation for robots to function as trusted teammates [10], rather than mere tools, within human-robot teams.
HARMONIC incorporates the cognitive architecture OntoAgent [11, 12, 9] as a strategic layer. Within HARMONIC, OntoAgent is responsible for semantic interpretation of sensory perception, grounding, attention management, goal-setting, sophisticated planning (supporting diverse multi-agent planning strategies), and addressing unexpected challenges in ways that humans can readily interpret. Additionally, OntoAgent supports meaning-oriented natural language communication, enabling clear causal explanations for the robots’ actions and decisions [10]. This capability significantly enhances skill transfer and knowledge sharing within human-robot teams. The use of the OntoAgent framework greatly enhances transparency of decision-making in HARMONIC-based robots, establishing crucial foundations for trust in human-robot collaboration [9].
@@
We present our approach to human-robot teaming using the example of a team comprised of two robots and a human executing a joint task. We demonstrate the robots’ knowledge of team organization as well as their ability to (a) dynamically negotiate tasks and allocate responsibilities, (b) select and modify plans, and (c) adapt language communication to both the situation and the needs of their interlocutors. We implement these robotic capabilities using the novel HARMONIC architecture that operates in parallel at two levels – the strategic cognitive level, which supports operations requiring reasoning, and the tactical control level, which supports skill-based, “automatic,” reflexive operation of robots.
Specifically, the key contributions of this paper are as follows:
@@
  • We introduce HARMONIC, a cognitive-robotic architecture, which enables a flexible integration of OntoAgent [11], a cognitive architecture with robot planning and control using Behavior Trees [13].
@@
  • Results of simulation experiments on a search task carried out by a robotic team, which showcases the human-robot team’s interactions and ability to handle complex, scenarios.
Section II briefly describes the state of the art and background. Section III introduces the HARMONIC framework and the implementation of its strategic and tactical levels. Section IV describes the task, the team configuration, and the development of distributed plans, and Section V presents the evaluation.
@@ II-A Cognitive-Robotic Architectures
There are relatively few cognitive architectures explicitly designed or effectively adapted for robotic applications. Among these architectures, Soar [14], ACT-R [15], and DIARC [16] are prominent, each contributing distinct methods and philosophies toward cognitive autonomy in robotics.
@@
DIARC is explicitly designed for cognitive robotics with distributed modules [16, 17], affective processes, and natural language capabilities suitable for dynamic human-robot interactions in narrow domains. However, it struggles with concurrent process management and lacks clear architectural layering, unified ontology, explainability, long-term episodic memory, explicit metacognition, and formalized safety mechanisms, thus limiting its robustness and adaptability. In summary, most architectures require extensive customization from both cognitive and control perspectives, leading to limited real-world applications and practical testing.
@@ II-B Multi-Agent Planning (MAP) and Communication
Multi-Agent Planning (MAP) in Heterogeneous Multi-Robot (HMR) teams includes a variety of centralized, decentralized, and hybrid strategies[20, 5, 21], each presenting distinct trade-offs. None of the current MAP approaches incorporates sufficient means of intra-team communication, which is an essential requirement for robust HMR teaming. Such communication must involve distinct protocols for human-robot and robot-robot interactions [24]. A unified, human-centric communication approach within a common cognitive architecture is advocated by Lemon et al. [25] and Natarajan et al. [7]. HARMONIC, with OntoAgent as its cognitive strategic layer, implements human-oriented communication, reasoning, and team dynamics. It also incorporates a variety of these scientific advances, such as the notion of common ground introduced by Clark et al. [26] and further developed by Klein et al. [27], the latest in dialogue processing systems [28], grounding techniques [29, 30], work on team organization [31] and natural language communication [9].
@@ II-C The OntoAgent Cognitive Architecture
OntoAgent is a cognitive architecture, designed to support the development of social intelligent agents supporting content-centric computational cognitive modeling [12, 9, 11]. This approach emphasizes the need to acquire, maintain, and dynamically expand large-scale knowledge bases, which are essential for an agent’s perception interpretation, reasoning, and action. The architecture’s memory structure is divided into three main components: a Situation Model (SM) that contains currently active concept instances; a Long-Term Semantic Memory (LTS) that stores knowledge about instances of events and objects, and an Episodic Memory (LTE) that stores knowledge about instances of events and objects. OntoAgent supports goal-directed behavior through a goal agenda and a prioritizer that selects which goals to pursue. It typically uses stored plans associated with goals but can also engage in reasoning from first principles when necessary.
The architecture is built on a service-based infrastructure, with key services including perception interpretation, attention management, reasoning, and action rendering. OntoAgent can integrate both native services and imported external capabilities, such as robotic vision or text-to-speech modules. This flexibility allows it to support multiple perception channels, including language understanding, visual perception, and simulated interoception (Chapter 8 of [12]).
@@ III The HARMONIC Framework
HARMONIC (Human-AI Robotic Team Member Operating with Natural Intelligence and Communication), shown in Fig. 1, is a dual-control cognitive-robotic architecture that integrates strategic, cognitive-level decision-making with tactical, skill-level robot control. It builds upon and extends the concept of hybrid control systems and architectures, as discussed in a comprehensive review by Dennis et al. [32]. It also represents an advance over the type 2 integration approach employed in DIARC [16, 17]. Unlike DIARC, where the strategic layer is embedded as a subsystem within the tactical layer to enable concurrent and dynamic operations, HARMONIC introduces a more sophisticated integration of these components.
@@
When the strategic layer receives perception inputs, including speech, vision, and robot state information, as data frames from the tactical layer, it interprets them in the context of its active goals and situation model. These inputs are transformed into ontologically grounded meaning representations, enabling unified reasoning across modalities. This interpretation supports downstream processes such as attention management and strategic decision-making by allowing the agent to evaluate the relevance of new information, update its understanding of the environment, monitor progress on current plans, and determine whether adjustments are necessary. Based on this reasoning, the OntoAgent in the strategic layer may issue updated action commands or generate natural language responses to teammates.
The tactical layer, on the other hand, is responsible for reflexive attention management, robot action planning and control, sensor input processing, and executing motor actions based on high-level commands received from the strategic layer. The attention manager processes perception data, translating and prioritizing sensory information for use by the strategic component. Additionally, the tactical layer employs specialized controllers, algorithms, and models to convert abstract commands from OntoAgent in the strategic layer into precise, actionable robot operations.
For example, a command such as PICK(key, at-position1), from the OntoAgent to “pick up a key on the ground,” triggers a well-defined sequence of tasks and actions, including object identification, positional assessment, trajectory computation, and actuation, in the tactical layer. Furthermore, the tactical layer manages reactive behaviors, including collision avoidance, through dedicated controllers.
The tactical layer of HARMONIC is built around the Behavior Trees (BTs) [13], which enable effective reactive control and support the system’s safety and operational requirements in dynamic environments. The tactical layer may also include advanced control policies for whole-body compliant control [33, 13] with motion planning, path planning [34], etc. BTs can also be used for representing skills and low-level plans that can also be shared between the robots [35, 36].
In the context of HARMONIC, the state manager module (a blackboard) keeps track of condition and state variables on the tactical layer [37]. As shown in Fig. 2, this allows for efficient querying and updating of the system’s state during operation based on the sensory inputs and action commands received from the strategic layer. Sensory inputs are continuously written into the State Manager (SM), where the attention module filters and packages this perception data for the OntoAgent in the strategic layer. Simultaneously, the high-level action commands sent by the OntoAgent through the Interface APIs are also unpacked for updating the variables in the SM.
While sensor data and physical state tracking operate in real time, the strategic layer runs asynchronously. Real-time control is achieved through the data available in the SM, which is accessed in parallel by the LiveBT (a BT instance), a module that chooses policies and actions stored in the Action Manager libraries. This architecture enables HARMONIC to maintain real-time behavior while supporting higher-level planning capabilities.
@@
These APIs update the placeholder variables within the State Manager, ensuring accurate and timely control of the tactical layer operations.
The HARMONIC architecture aligns with Kahneman’s dual-system approach [38], with the OntoAgent in the strategic layer implementing System 2 (slow, deliberative reasoning) and the BTs in the tactical layer implementing System 1 (fast, intuitive responses). This structure allows for integrating reflexive and deliberative functioning, flexible scheduling and adaptation, which in turn enables real-time prioritization of goals and actions. As a result, HARMONIC effectively handles computational delays, contingencies, safety concerns, and resource optimization through a combination of low-level planning and reactive algorithms.
@@ IV Task Design
We evaluated the HARMONIC framework’s teaming strategy using a simulated search task set in an apartment environment, as illustrated in Fig. 4.6. The scenario involves an HMR team consisting of an unmanned ground vehicle (UGV), a drone, and a human team member. The objective is to find a set of lost keys. The human initiates the task by remotely communicating with the robotic team members in English. The simulation environment, developed in Unity, incorporates abstracted versions of robot behaviors. This approach allowed us to focus on key aspects of the HARMONIC framework’s performance while simplifying the implementation of complex robot actions that are not central to our current work.
@@ IV-A Heterogenous Robot Team Configuration
The simulation evaluation involved two functionally and structurally heterogeneous robots – a UGV and a drone – offering complementary capabilities and perspectives [39]. Each robot is equipped with sensors for localization, collision detection, and object recognition, facilitating coordinated and efficient operations.
The UGV’s (see Fig. 3.a) can access ground-level areas and maneuver under obstacles and surfaces. The drone, on the other hand, complements the UGV by performing aerial search operations. The drone’s (see Fig. 3.b) vertical mobility and sensing capabilities allow it to scan environments extensively from above, navigating over obstacles and accessing areas difficult for the UGV to access. Since the robots communicate in natural language, their dialogue and actions are comprehensible to humans, and humans can intervene when necessary.
Both robots operate independent instances of HARMONIC, with different content parts (for example, there is no presumption that robots have identical ontological world models). For the experiment, the UGV acts as the team leader at the task level, responsible for managing interactions with the human operator and coordinating with the drone. The team leader assignment is presently arbitrary; however, future work aims to develop and evaluate heuristic methods for dynamic team hierarchies. This advancement will optimize leader selection based on contextual factors and the individual capabilities of each agent relevant to specific tasks.